OpenStudy (tia_sweetheart17):

what is 3x^4 y^2 over x y^2

OpenStudy (anonymous):

\(\huge \frac{3x^4 y^2}{x y^2}=\frac{3x^{\cancel{4}^3} \cancel{y^2}}{\cancel{x} \cancel{y^2}}= \) ????